Find the greatest possible pair of integers such that one integer is twice the other and their sum is less than or equal to 12.
Answer by Theo(13342)

x = one of the integers
y = the other

x = 2*y
x + y <= 12

substitute for x in the second equation to get:

2*y + y <= 12

combine like terms to get:

3*y <= 12

divide both sides by 3 to get:

y <= 4

the greatest that y can be is 4.

if y is equal to 4, then x = 2*y means that x = 8

the greatest that x can be is 8.

when x = 8 and y = 4, then:

8 = 2*4 = ok
4 + 8 = 12 = ok